# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
$\lim_{k\to\infty}\int_{B_{R}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}^{\ast}(x)|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x=\infty.$ (3.21)
@@
$-\Delta u_{k}=g_{k}\quad\text{in}\quad\mathbb{R}^{n}\quad\text{and}\quad-\Delta u=g\quad\text{in}\quad\mathbb{R}^{n}.$
It is obvious that $g_{k},g\in L^{t}(\mathbb{R}^{n})$ for all $k\in\mathbb{N}$ and for any $t\in\{p,\frac{n}{2}\}$. Define $g_{k}^{\ast}(x)=g_{k}^{\sharp}(\sigma_{n}|x|^{n})$ for all $x\in\mathbb{R}^{n}$ and $k\in\mathbb{N}$. Hence we obtain from (3.17) that $\|g_{k}\|_{t}\leq 1$ for all $k\in\mathbb{N}$ and $g_{k}\rightharpoonup g$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$. Now by using Lemma 2.2, there exists a subsequence of $\{g_{k}\}_{k}$ still not relabelled and $h\in L^{t}([0,+\infty))$ such that
$g_{k}^{\sharp}\to h~{}\text{a.e. in}~{}[0,+\infty)~{}~{}\text{as}~{}~{}k\to\infty~{}~{}\text{and}~{}~{}\|h\|_{L^{t}([0,+\infty))}\geq\|g^{\sharp}\|_{L^{t}([0,+\infty))}=\|g\|_{t}~{}~{}~{}\text{for any }~{}t\in\{p,\frac{n}{2}\}.$ (3.22)
@@
$\quad\quad u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})$ $\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+D(n)\|g_{k}\|_{\frac{n}{2}}\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\bigg{(}\int_{0}^{t}g_{k}^{\sharp}(\eta)\,\ \mathrm{d}\eta\bigg{)}t^{\frac{2}{n}-2}\,\mathrm{d}t+D(n)$
@@
Due to Hölder’s inequality and (3.22), we obtain for all $s\in(0,+\infty)$ that
@@
Moreover, we deduce from the above inequalities that there exists a suitable constant $\overset{\sim}{D}(n)>0$ which is depending on $n$ but independent on the choice of $k$ and there holds
$u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\ \mathrm{d}t+\overset{\sim}{D}(n),~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad 0<s_{1}<s_{2}<+\infty.$ (3.23)
@@
$u_{k}^{\sharp}(s)$ $\leq w_{k}(s)+u_{k}^{\sharp}(|B_{R}|)+\overset{\sim}{D}(n)\leq w_{k}(s)+\overset{\sim}{D}(n,R),~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|),$
[…]
$|u_{k}^{\sharp}(s)|^{\frac{n}{n-2}}\leq(1+\eta)|w_{k}(s)|^{\frac{n}{n-2}}+C_{\eta}\overset{\sim}{D}(n,R)^{\frac{n}{n-2}},~{}\forall~{}k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|).$
@@
$\leq\operatorname{exp}(\ell\beta_{\gamma,n}C_{\eta}\overset{\sim}{D}(n,R)^{\frac{n}{n-2}})\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s.$
@@
$\lim_{k\to\infty}\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s=\infty.$ (3.24)
In virtue of Hölder’s inequality and (3.22), for all $s\in(0,|B_{R}|)$, we have
$w_{k}(s)$ $\leq\bigg{(}\int_{0}^{|B_{R}|}|g_{k}^{\sharp}(t)|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\int_{s}^{|B_{R}|}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg{)}^{1-\frac{2}{n}}$
$\leq\|g^{\ast}_{k}\|_{\frac{n}{2}}\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}}\leq\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}},~{}\forall~{}k\in\mathbb{N}.$ (3.25)
Next, we claim that for any (¯$\ell^{\prime}\in(\bar{\ell},L_{n}(u),~{}k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$, there exist $k\in\mathbb{N}$ and $s\in(0,s_{0})$ satisfying
@@
Indeed, if not, let there exists $k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$ be such that for some (¯$\ell^{\prime}\in(\bar{\ell},L_{n}(u)$, we have
$w_{k}(s)<\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s}\bigg{)}\bigg{)}^{1-\frac{2}{n}},~{}\forall~{}k\geq k_{0}\quad\text{and}\quad s\in(0,s_{0}).$ (3.26)
By using (3) and (3.26), for sufficiently large $k$, we have
$\int_{0}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})$ $\,\mathrm{d}s=\bigg{(}\int_{0}^{s_{0}}+\int_{s_{0}}^{|B_{R}|}\bigg{)}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s$
$\leq\int_{0}^{s_{0}}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\bigg{(}\frac{|B_{R}|}{s}\bigg{)}^{(1-\frac{\gamma}{n})\frac{\bar{\ell}}{\ell^{\prime}}}\,\mathrm{d}s+\int_{s_{0}}^{|B_{R}|}\Big{(}\frac{\sigma_{n}}{s}\Big{)}^{\frac{\gamma}{n}}\bigg{(}\frac{|B_{R}|}{s}\bigg{)}^{(1-\frac{\gamma}{n})\bar{\ell}}\,\mathrm{d}s$
<+<$\infty$.< + ∞ .
In view of the above inequality and (3.24), we arrive at a contradiction. This completes the proof of our claim. Thus there exists a subsequence $\{s_{k}\}_{k}\subset(0,s_{0})$ with $s_{k}\leq\frac{1}{k}$ for all $k\in\mathbb{N}$ such that we can assume
$w_{k}(s_{k})\geq\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}.$ (3.27)
@@
In view of (3.27), one can easily see that $w_{k}(s_{k})\nearrow\infty$ as $k\to\infty$ and also we have $w_{k}(|B_{R}|)=0$. This shows that we can find a subsequence of $\{s_{k}\}_{k}$ still not relabelled and the existence of $\xi_{k}\in(s_{k},|B_{R}|)$ satisfying $w_{k}(\xi_{k})=L$ for each $k\in\mathbb{N}$. In addition, since the map $\cdot\mapsto g_{k}^{\sharp}(\cdot)$ is always nonnegative and nonincreasing, therefore by using Hölder’s inequality, one has
@@
$\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}-L\leq w_{k}(s_{k})-w_{k}(\xi_{k})\leq\frac{ng_{k}^{\sharp}(s_{k})}{2\beta(n,2)^{1-\frac{2}{n}}}\Big{(}|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big{)},~{}\forall~{}k\in\mathbb{N}.$ (3.28)
Now from (3.28), we infer that $g_{k}^{\sharp}(s_{k})\nearrow\infty$ as $k\to\infty$ and thus there exists $k_{0}\in\mathbb{N}$ such that $g_{k}^{\sharp}(s_{k})>L$ for all $k\geq k_{0}$. Consequently, up to a subsequence of $\{s_{k}\}_{k}$ still not relabelled, we can find $\eta_{k}\in(s_{k},|B_{R}|)$ satisfying $g_{k}^{\sharp}(\eta_{k})=L$ for each $k\in\mathbb{N}$ and $g_{k}^{\sharp}(\eta)<L$ for all $\eta>\eta_{k}$. For all $k\geq k_{0}$, we set $\theta_{k}:=\min\{\eta_{k},\xi_{k}\}$, then one sees from (3.27) that
$\bigg{(}\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}-L$ $\leq w_{k}(s_{k})-w_{k}(\xi_{k})=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\xi_{k}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$
@@
$\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\Bigg{[}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+\frac{nL}{2}\Big{(}|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big{)}\Bigg{]}.$ (3.29)
On the other hand, the application of Hölder’s inequality yields
$\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$ $\leq\bigg{(}\int_{s_{k}}^{\theta_{k}}|g_{k}^{\sharp}(t)-L|^{\frac{n}{2}}\,\mathrm{d}t\bigg{)}^{\frac{2}{n}}\bigg{(}\int_{s_{k}}^{\theta_{k}}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg{)}^{1-\frac{2}{n}}$
@@
$\leq\|H_{L}(g^{\sharp}_{k})\|_{L^{\frac{n}{2}}([0,+\infty))}\bigg{(}\frac{1}{\beta(n,2)}\ln\bigg{(}\frac{|B_{R}|}{s_{k}}\bigg{)}\bigg{)}^{1-\frac{2}{n}}.$ (3.30)
If possible, let us choose $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$. It is obvious that $\ln\big{(}\frac{|B_{R}|}{s_{k}}\big{)}\to\infty$ as $k\to\infty$. Hence we obtain from (3) and (3) as $k\to\infty$ that
@@
$\leq\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$ (3.31)
@@
11 11 $=\|u_{k}\|^{\frac{n}{2}}=\|g_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g_{k}\|^{\frac{n}{2}}_{p}$
@@
$\quad+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$ (3.32)
In view of (3) and (3), we get as $k\to\infty$ that
1($1-\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}\geq\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$ (3.33)
@@
$h(s)-H_{L}(h(s))\to h(s)\quad\text{as}\quad L\to\infty\quad\text{and}\quad g^{\sharp}_{k}(s)-H_{L}(g^{\sharp}_{k}(s))\to h(s)-H_{L}(h(s))\quad\text{as}\quad k\to\infty.$
@@
1($1-\bigg{(}\frac{1}{\overset{\sim}{\ell}}\bigg{)}^{\frac{n-2}{2}}\geq\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$
@@
which contradicts $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$ and thus we complete the proof of the first part of Theorem 1.4.
To complete the proof, we only have to prove the sharpness of $L_{n}(u)$. For this, we shall have to contruct a sequence $\{u_{k}\}_{k}\subset E$ and a function $u\in E$ satisfying
$\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty,\quad\|u\|=\delta<1$ (3.34)
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow\infty\quad\text{as}\quad k\to\infty,\quad\forall~{}\ell>L_{n}(u).$
Let $\{\xi_{k}\}_{k}\subset E$ be a sequence of radial functions, as defined in (3.4). It follows from the standard arguments that
@@
Take a radial function $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u\|=\delta<1$ and $\text{supp}(u)\subset\{x\in\mathbb{R}^{n}:~{}2<|x|<3\}$. Define $v_{k}:=u+(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\xi_{k}$ for all $k\in\mathbb{N}$. Notice that $\Delta u$ and $\Delta\xi_{k}$ have disjoint supports. It follows at once that
@@
where $\zeta_{k}\to 0$ as $k\to\infty$. Consequently, one sees that $\|v_{k}\|=1+\zeta_{k}$. Denote $u_{k}:=\frac{v_{k}}{1+\zeta_{k}}$. This implies the validity of (3.34). Moreover, for any $\ell>L_{n}(u)$, there exists $\varepsilon>0$ such that we can assume $\ell=(1+\varepsilon)L_{n}(u)$. In conclusion, we have the following esimates

### v3

@@ 2. Preliminary results
(4.15) $\lim_{k\to\infty}\int_{B_{R}}\frac{\operatorname{\exp}(\ell\beta_{\gamma,n}|u_{k}^{\ast}(x)|^{\frac{n}{n-2}})}{|x|^{\gamma}}\,\mathrm{d}x=+\infty.$
@@
$\Delta u_{k}=g_{k}\quad\text{in}\quad\mathbb{R}^{n}\quad\text{and}\quad-\Delta u=g\quad\text{in}\quad\mathbb{R}^{n}.$
It is obvious that $g_{k},g\in L^{t}(\mathbb{R}^{n})$ for all $k\in\mathbb{N}$ and for any $t\in\{p,\frac{n}{2}\}$. Define $g_{k}^{\ast}(x)=g_{k}^{\sharp}(\sigma_{n}|x|^{n})$ for all $x\in\mathbb{R}^{n}$ and $k\in\mathbb{N}$. Hence, we obtain from (4.11) that $\|g_{k}\|_{t}\leq 1$ for all $k\in\mathbb{N}$ and $g_{k}\rightharpoonup g$ in $L^{t}(\mathbb{R}^{n})$ as $k\to\infty$ for any $t\in\{p,\frac{n}{2}\}$. Now, by using Lemma 2.2, there exists a subsequence of $\{g_{k}\}_{k}$ still not relabeled and $h\in L^{t}([0,+\infty))$ such that for any $t\in\{p,\frac{n}{2}\}$, we have
(4.16) $g_{k}^{\sharp}\to h~\text{a.e. in}~[0,+\infty)~~\text{as}~~k\to\infty~~\text{and}~~\|h\|_{L^{t}([0,+\infty))}\geq\|g^{\sharp}\|_{L^{t}([0,+\infty))}=\|g\|_{t}.$
@@
$\quad\quad u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})$ $\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+D(n)\|g_{k}\|_{\frac{n}{2}}$
$\leq\frac{1}{n^{2}\sigma_{n}^{\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\bigg(\int_{0}^{t}g_{k}^{\sharp}(\eta)\,\ \mathrm{d}\eta\bigg)t^{\frac{2}{n}-2}\,\mathrm{d}t+D(n)$
@@
Due to the Hölder’s inequality and (4.16), we obtain for all $s\in(0,+\infty)$ that
@@
Moreover, we deduce from the above inequalities that there exists a suitable constant $\overset{\sim}{D}(n)>0$ which depends on $n$ but is independent of the choice of $k$ and there holds
(4.17) $u_{k}^{\sharp}(s_{1})-u_{k}^{\sharp}(s_{2})\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{1}}^{s_{2}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\ \mathrm{d}t+\overset{\sim}{D}(n),~\forall~k\in\mathbb{N}\quad\text{and}\quad 0<s_{1}<s_{2}<+\infty.$
@@
$u_{k}^{\sharp}(s)$ $\leq w_{k}(s)+u_{k}^{\sharp}(|B_{R}|)+\overset{\sim}{D}(n)\leq w_{k}(s)+\overset{\sim}{D}(p,n,R),~\forall~k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|),$
where $\overset{\sim}{D}(p,n,R)>0$ is a constant depending only on $p$, $n$ and $R$. Choose $\eta>0$ sufficiently small enough such that $\bar{\ell}=(1+\eta)\ell<L_{n}(u)$, then from the above inequality together with (3.6) yields
$|u_{k}^{\sharp}(s)|^{\frac{n}{n-2}}\leq(1+\eta)|w_{k}(s)|^{\frac{n}{n-2}}+C_{\eta}\overset{\sim}{D}(p,n,R)^{\frac{n}{n-2}},~\forall~k\in\mathbb{N}\quad\text{and}\quad s\in(0,|B_{R}|).$
@@
$\leq\operatorname{exp}(\ell\beta_{\gamma,n}C_{\eta}\overset{\sim}{D}(p,n,R)^{\frac{n}{n-2}})\int_{0}^{|B_{R}|}\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s.$
@@
(4.18) $\lim_{k\to\infty}\int_{0}^{|B_{R}|}\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s=+\infty.$
In virtue of the Hölder’s inequality and (4.16), for all $s\in(0,|B_{R}|)$, we have
(4.19) $w_{k}(s)$ $\leq\bigg(\int_{0}^{|B_{R}|}|g_{k}^{\sharp}(t)|^{\frac{n}{2}}\,\mathrm{d}t\bigg)^{\frac{2}{n}}\bigg(\int_{s}^{|B_{R}|}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg)^{1-\frac{2}{n}}$
$\leq\|g^{\ast}_{k}\|_{\frac{n}{2}}\bigg(\frac{1}{\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s}\bigg)\bigg)^{1-\frac{2}{n}}\leq\bigg(\frac{1}{\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s}\bigg)\bigg)^{1-\frac{2}{n}},~\forall~k\in\mathbb{N}.$
Next, we claim that for any $\ell^{\prime}\in(\bar{\ell},L_{n}(u)),~k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$, there exist $k\in\mathbb{N}$ and $s\in(0,s_{0})$ satisfying
@@
Indeed, if not, let there exists $k_{0}\in\mathbb{N}$ and $s_{0}\in(0,|B_{R}|)$ be such that for some $\ell^{\prime}\in(\bar{\ell},L_{n}(u))$, we have
(4.20) $w_{k}(s)<\bigg(\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s}\bigg)\bigg)^{1-\frac{2}{n}},~\forall~k\geq k_{0}\quad\text{and}\quad s\in(0,s_{0}).$
By using (4.19) and (4.20), for sufficiently large $k$, we have
$\int_{0}^{|B_{R}|}\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}$ $\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s$
$=\bigg(\int_{0}^{s_{0}}+\int_{s_{0}}^{|B_{R}|}\bigg)\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}\operatorname{\exp}(\bar{\ell}\beta_{\gamma,n}|w_{k}(s)|^{\frac{n}{n-2}})\,\mathrm{d}s$
$\leq\int_{0}^{s_{0}}\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}\bigg(\frac{|B_{R}|}{s}\bigg)^{(1-\frac{\gamma}{n})\frac{\bar{\ell}}{\ell^{\prime}}}\,\mathrm{d}s+\int_{s_{0}}^{|B_{R}|}\Big(\frac{\sigma_{n}}{s}\Big)^{\frac{\gamma}{n}}\bigg(\frac{|B_{R}|}{s}\bigg)^{(1-\frac{\gamma}{n})\bar{\ell}}\,\mathrm{d}s<+\infty.$
In view of the above inequality and (4.18), we arrive at a contradiction. This completes the proof of our claim. Thus, there exists a subsequence $\{s_{k}\}_{k}\subset(0,s_{0})$ with $s_{k}\leq\frac{1}{k}$ for all $k\in\mathbb{N}$ such that we can assume
(4.21) $w_{k}(s_{k})\geq\bigg(\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s_{k}}\bigg)\bigg)^{1-\frac{2}{n}}.$
@@
In view of (4.21), one can easily see that $w_{k}(s_{k})\nearrow+\infty$ as $k\to\infty$ and also we have $w_{k}(|B_{R}|)=0$. This shows that we can find a subsequence of $\{s_{k}\}_{k}$ still not relabeled and the existence of $\xi_{k}\in(s_{k},|B_{R}|)$ satisfying $w_{k}(\xi_{k})=L$ for each $k\in\mathbb{N}$. In addition, since the map $\cdot\mapsto g_{k}^{\sharp}(\cdot)$ is always nonnegative and nonincreasing, therefore by using the Hölder’s inequality, one has
@@
(4.22) $\bigg(\frac{1}{\ell^{\prime}\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s_{k}}\bigg)\bigg)^{1-\frac{2}{n}}-L\leq w_{k}(s_{k})-w_{k}(\xi_{k})\leq\frac{ng_{k}^{\sharp}(s_{k})}{2\beta(n,2)^{1-\frac{2}{n}}}\Big(|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big),~\forall~k\in\mathbb{N}.$
Now, from (4.22), we infer that $g_{k}^{\sharp}(s_{k})\nearrow+\infty$ as $k\to\infty$ and thus, there exists $k_{0}\in\mathbb{N}$ such that $g_{k}^{\sharp}(s_{k})>L$ for all $k\geq k_{0}$. Consequently, up to a subsequence of $\{s_{k}\}_{k}$ still not relabeled, we can find $\eta_{k}\in(s_{k},|B_{R}|)$ satisfying $g_{k}^{\sharp}(\eta_{k})=L$ for each $k\in\mathbb{N}$ and $g_{k}^{\sharp}(\eta)<L$ for all $\eta>\eta_{k}$. For all $k\geq k_{0}$, we set $\theta_{k}:=\min\{\eta_{k},\xi_{k}\}$, then one sees from (4.21) that
(4.23) $\bigg(\frac{1}{\ell^{\prime}\beta(n,2)}$ $\ln\bigg(\frac{|B_{R}|}{s_{k}}\bigg)\bigg)^{1-\frac{2}{n}}-L$
$\leq w_{k}(s_{k})-w_{k}(\xi_{k})=\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\xi_{k}}\frac{g_{k}^{\sharp}(t)}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$
@@
$\leq\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\Bigg[\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t+\frac{nL}{2}\Big(|B_{R}|^{\frac{2}{n}}-s_{k}^{\frac{2}{n}}\Big)\Bigg].$
Moreover, the application of the Hölder’s inequality yields
(4.24) $\frac{1}{\beta(n,2)^{1-\frac{2}{n}}}\int_{s_{k}}^{\theta_{k}}\frac{g_{k}^{\sharp}(t)-L}{t^{1-\frac{2}{n}}}\,\mathrm{d}t$ $\leq\bigg(\int_{s_{k}}^{\theta_{k}}|g_{k}^{\sharp}(t)-L|^{\frac{n}{2}}\,\mathrm{d}t\bigg)^{\frac{2}{n}}\bigg(\int_{s_{k}}^{\theta_{k}}\frac{1}{\beta(n,2)t}\,\mathrm{d}t\bigg)^{1-\frac{2}{n}}$
@@
$\leq\|H_{L}(g^{\sharp}_{k})\|_{L^{\frac{n}{2}}([0,+\infty))}\bigg(\frac{1}{\beta(n,2)}\ln\bigg(\frac{|B_{R}|}{s_{k}}\bigg)\bigg)^{1-\frac{2}{n}}.$
If possible, let us choose $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$. It is obvious that $\ln\Big(\frac{|B_{R}|}{s_{k}}\Big)\to+\infty$ as $k\to\infty$. Hence, we obtain from (4.23) and (4.24) as $k\to\infty$ that
@@
$\leq\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$
@@
(4.25) $1$ $=\|u_{k}\|^{\frac{n}{2}}=\|g_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}+\|g_{k}\|^{\frac{n}{2}}_{p}$
@@
$\quad+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$
In view of (4) and (4.25), we get as $k\to\infty$ that
(4.26) $1-\Bigg(\frac{1}{\overset{\sim}{\ell}}\Bigg)^{\frac{n-2}{2}}\geq\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|g^{\sharp}_{k}-H_{L}(g^{\sharp}_{k})\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}+o_{k}(1).$
@@
$h(s)-H_{L}(h(s))\to h(s)\quad\text{as}\quad L\to\infty$
and
$g^{\sharp}_{k}(s)-H_{L}(g^{\sharp}_{k}(s))\to h(s)-H_{L}(h(s))\quad\text{as}\quad k\to\infty.$
@@
$1-\Bigg(\frac{1}{\overset{\sim}{\ell}}\Bigg)^{\frac{n-2}{2}}\geq\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{\frac{n}{2}}([0,+\infty))}+\|h-H_{L}(h)\|^{\frac{n}{2}}_{L^{p}([0,+\infty))}.$
@@
which contradicts $\overset{\sim}{\ell}\in(\ell^{\prime},L_{n}(u))$ and thus, we complete the proof of the first part of Theorem 1.2. To complete the proof, we only have to prove the sharpness of $L_{n}(u)$. For this, we shall have to contruct a sequence $\{u_{k}\}_{k}\subset E$ and a function $u\in E$ satisfying
(4.27) $\|u_{k}\|=1,\quad u_{k}\rightharpoonup u\quad\text{in}\quad E\quad\text{as}\quad k\to\infty,\quad\|u\|=\delta<1$
@@
$\int_{\mathbb{R}^{n}}\frac{\Phi_{\ell\beta_{\gamma,n},j_{0}}(u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\nearrow+\infty\quad\text{as}\quad k\to\infty,~\forall~\ell>L_{n}(u).$
Let $\{\xi_{k}\}_{k}\subset E$ be a sequence of radial functions defined as in (3.4). It follows from the standard arguments that
@@
Take a radial function $u\in C^{\infty}_{0}(\mathbb{R}^{n})$ such that $\|u\|:=\delta<1$ and $\text{supp}(u)\subset\{x\in\mathbb{R}^{n}:~2<|x|<3\}$. Define $v_{k}:=u+(1-\delta^{\frac{n}{2}})^{\frac{2}{n}}\xi_{k}$ for all $k\in\mathbb{N}$. Notice that $\Delta u$ and $\Delta\xi_{k}$ have disjoint supports. It follows at once that
@@
where $\zeta_{k}\to 0$ as $k\to\infty$. Hence, it is easy to see that $\|v_{k}\|=1+\zeta_{k}$. Denote $u_{k}:=\frac{v_{k}}{1+\zeta_{k}}$. This implies the validity of (4.27). Moreover, for any $\ell>L_{n}(u)$, there exists $\varepsilon>0$ such that one can assume $\ell=(1+\varepsilon)L_{n}(u)$ and thus, we have